# Measurement of the dynamic charge susceptibility near the charge density wave transition in ErTe3

### v1

@@ Abstract
A key thermodynamic feature of a CDW is traditionally believed to be a divergence in the static charge susceptibility, $\chi(q)$, which favors a periodic redistribution of charge at low temperatures [1]. This redistribution causes a gap to open in the single-particle density of states and is accompanied by a periodic distortion of the structural lattice, driven by electron-phonon coupling—a phenomenon known as a Peierls instability. While this divergence is most pronounced in one-dimensional systems, where Fermi surface nesting may be important [5], CDW instabilities have abeen widely observed in higher-dimensional systems. In these cases, nesting is imperfect and strongly dependent on the specific morphology of the Fermi surface [7, 8].
The precise conditions under which the charge susceptibility exhibits a divergence, however, are not well understood. The nesting picture by itself is usually too simplistic, and effects such as temperature, scattering, and band velocity mismatch at the nesting wave vector play a crucial role in real materials [8, 9]. Coincidence between the nesting vector and the CDW wave vector is rarely observed in nature, and other aspects such as the momentum dependence of the electron-phonon interaction have been invoked to quantitatively relate theory and experiments [10, 7, 11, 12]. It is therefore crucial to experimentally measure the low-energy charge dynamics of a prototypical CDW material to understand the collective excitations that drive the CDW instability, and to determine whether, and how, a susceptibility divergence occurs.
@@
Figure 1: (a) Crystal structure of ErTe3 showing the corrugated Er-Te slabs and Te square nets. (b) Schematic of the Fermi surface showing the nesting vectors (adapted from Ref. [43]). (c) Scattering geometry of the experiment with reference to the sample orientation. ki and ks represent momenta of the incident and scattered electron, respectively, and $q$ denotes the in-plane momentum transfer which is along $c$-axis.
@@
Figure 2: Elastic response: (a) m-EELS spectra ($\omega=0$) as a function of $q_{\|}$ for different temperatures. (b) The integrated area under the CDW peak $q\sim q_{0}$ obtained from fits to Eq.(2) along with the fit to mean-field interpolation formula, normalized to the value in the $T\rightarrow 0$. (c) $w(t)=\gamma_{CDW}/\gamma_{(0,1)}$, the ratio of the width of the CDW satellite to that of the Bragg peak as a function of temperature. (d) Ordering wave vector $q_{0}$ as a function of temperature. Vertical gray dashed line in (c,d) denotes the Tc obtained in fit in (b).
@@
where $G(Q,T)=1/[\kappa Q^{2}+\mu(T)]$ is a Lorentzian, $\sigma$ represents the disorder strength, $Q=q-q_{0}$ is the momentum relative to the CDW wave vector, and the other parameters quantify the CDW stiffness and correlation length. The first term in Eq. (1) is truly static, while the second is dynamical and, in principle, could be eliminated by using an energy analyzer that perfectly isolated the elastic ($\omega=0$) response.
@@
where $q_{0}(T)$ is the CDW ordering wavevector and the width, $\gamma_{q}(T)$, is proportional to the inverse correlation length (see Supplementary Material). The sum $m(T)q+c(T)$ represents a linear background whose meaning will be discussed further below. Surprisingly, adding an additional, Lorentzian-squared term did not improve the quality of the fits. This is surprising given that our measurement is energy-resolved, with an overall resolution of 5.6 meV, and should isolate the Lorentzian-squared term in Eq. (1). We conclude that the CDW in ErTe3 is influenced by disorder, but extremely weakly.
@@
Figure 3: Dynamical response: Low energy m-EELS spectra normalized to the maximum intensity for different temperatures for (a) $q=0.6r.l.u.$ and (b) $q\sim q_{0}$. (c) $\gamma(q,T)$ extracted by fitting the loss spectra to (LABEL:loss_fit). The dashed lines are a guide to the eye. (d) Temperature dependence of $\gamma(q,T)$ for $q=0.6r.l.u.$ and $q\sim q_{0}$. (e) $\tau^{-1}_{0}$ and (f) $D$ obtained from fitting $\gamma(q,T)$ near $q_{0}$ to (5).
@@
The integrated intensity of elastic CDW reflection, $I_{CDW}(T)$, is plotted in Fig. 2(b). $I_{CDW}(T)$ is proportional to the square of the order parameter, $\Delta(T)$, and a fit to the BCS interpolation formula $\Delta(T)=\Delta_{0}\tanh{1.74\sqrt{T_{c}/T-1}}$, yields a $T_{C}=268(7)K$ which is consistent with the previous x-ray diffraction measurements [18]. In the intermediate regime between 100-250K, the integrated intensity is smaller than the mean field prediction, which was also observed in the evolution of the single particle gap in ARPES [19]. Vestiges of the CDW order is also observed above $T_{c}$, up to $T=$290 K, presumably due to pinning on disorder [5].
Fig. 2(c) shows the ratio of the width of the CDW to that of the (0,1) Bragg peak, i.e., $w(t)=\gamma_{CDW}/\gamma_{(0,1)}$, as a function of temperature. Below, 225K the width of the CDW is nearly the same as that of the Bragg peak, indicating that it is limited by the instrumental resolution. Above 225K, the CDW peak broadens, indicating that the CDW correlation length gets shorter as the temperature is increased through $T_{c}$.
@@
The CDW wavevector, $q_{0}(T)$, shows a non-monotonic dependence on temperature (Fig. 2(d)). This behavior has also been observed in x-ray measurements of other tritellurides [18] and is a common characteristic of incommensurate CDWs. The exact value of $q_{0}$ often reflects a balance between factors such as the Fermi surface morphology, lattice pinning, the compressibility of uncondensed electrons, and coupling to hidden order parameters [50, 51, 49]. These behavior of $q_{0}(T)$ suggests that the CDW in ErTe3 is shaped by a combination of these effects.
@@
Having establishing the structural properties of the CDW through elastic scattering, we now focus on its collective dynamics. Figures 3(a,b) present the temperature dependence of the M-EELS spectra at $q=0.6$ r.l.u. and near the CDW ordering wavevector, $q=q_{0}$. Across all temperatures and momenta, the spectra are dominated by a quasi-elastic response, with no evidence of dispersive collective excitations. This finding contrasts sharply with IXS studies, which reveal a softening of phonon modes near the CDW wavevector above $T_{C1}$ [25, 26]. Since M-EELS is more sensitive to valence band electronic excitations, while IXS primarily detects phonons, this suggests that the electronic excitations in ErTe3 behave quite differently from its lattice excitations. This conclusion is further reinforced by the absence of phonon peaks in the low-energy optical conductivity of ErTe3 [22], an effect attributed to strong metallic screening.
@@
Despite the absence of well-defined, propagating collective modes, the dynamic charge susceptibility shows clear temperature dependence. A broadening of the quasi-elastic response is evident in the raw spectrum (Fig. 3(a,b)), suggesting that the CDW undergoes nontrivial, relaxational dynamics at an energy scale of $\sim k_{B}T$.
@@
where $n(\omega,T)$ is the Bose occupation factor and $V_{\text{eff}}(\omega,q)$ is the Coulomb matrix element [32]. This minimal model, with only two adjustable parameters ($A(q,T)$ and $\gamma(q,T)$) provides excellent fits to the M-EELS data ((see Supplementary Material)), confirming that the charge dynamics in ErTe3 are predominantly relaxational.
@@
$\gamma(q,T)=\tau^{-1}+D(T)\;(q-q_{0})^{2}$ (5)
where $\tau^{-1}$ represents pure dissipation and $D(T)$ is the diffusion constant. Figs. 3(e,f) show plots of $\tau^{-1}$ and $D(T)$, respectively. The dissipation rate, $\tau^{-1}$, varies quadratically with temperature, consistent with the behavior of a good Fermi liquid. Intriguingly, $D(T)$ peaks around $\sim$250 K, slightly below $T_{c}$. This suggests that CDW excitations propagate diffusively, with the diffusivity reaching its maximum just below $T_{c}$, when the CDW amplitude is large but still subject to significant thermal fluctuations.
@@
Finally, our analysis allows us to quantify the behavior of the static susceptibility, $\chi(q)$, which is traditionally expected to diverge near a CDW transition [1]. The static susceptibility is given by the real part of the complex susceptibility $\chi(q,\omega)$ evaluated at zero frequency. While this could be obtained via a Kramers-Kronig transformation of the data, we can also directly compute it from the real part associated with Eq. (3),
@@
Finally, we note that, while the largest changes in the static susceptibility occur at $q_{0}$, the susceptibility shows significant temperature dependence at all momenta measured, which appear as a linear background in Eq. (2). This indicates that a substantial fraction of the valence electrons—even those not directly involved in CDW ordering—contribute significantly to the overall change in charge susceptibility. What underlying physics would drive such fluctuations everywhere in momentum space is an open question.
@@ Acknowledgements.
We thank Steven A. Kivelson and Anisha G. Singh for helpful discussions. M-EELS measurements were supported by the Center for Quantum Sensing and Quantum Materials, an Energy Frontier Research Center funded by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award DE-SC0021238. QJ and IRF (crystal growth and characterization) were supported by DOE BES contract DE-AC02-76SF00515. TCC (surface preparation) was supported by DOE BES contract DE-FG02-07ER46383. EF (theory) was supported by National Science Foundation grant DMR 2225920. QJ was also partially supported by a Geballe Laboratory for Advanced Materials Postdoctoral Fellowship. P.A. gratefully acknowledges support from the EPiQS program of the Gordon and Betty Moore Foundation, grant GBMF9452.

### v3

@@ Abstract
An expected thermodynamic signature of the onset of a CDW phase is a divergence in static charge susceptibility $\chi(q)$, which favors a periodic redistribution of charge in the ordered state [1]. This redistribution causes a gap to open in the single-particle density of states and is accompanied by a periodic distortion of the structural lattice driven by electron-phonon coupling. Although this phenomenon is most pronounced in one-dimensional systems where a Peierls instability due to Fermi surface nesting leads to logarithmic divergence [5], CDW phases have been widely observed in higher-dimensional systems. In these cases, nesting is imperfect and strongly dependent on the specific morphology of the Fermi surface [7, 8].
The precise microscopic mechanism driving CDW mehavior in most materials remains elusive. The nesting picture by itself is usually too simplistic, and effects such as temperature, scattering, and band velocity mismatch at the nesting wave vector play a crucial role in real materials [8, 9]. Correspondence between the CDW wave vector and the nesting properties of the Fermi surface are often difficult to discern [10]. Other aspects, such as the momentum dependence of the electron-phonon interaction, have been invoked to quantitatively relate theory and experiments [11, 7, 12, 13]. It is therefore crucial to experimentally measure the low-energy charge dynamics of a prototypical CDW material to understand the collective excitations that drive the CDW instability, and to determine whether, and how, a susceptibility divergence occurs.
@@
Figure 1: (a) Crystal structure of ErTe3 showing the corrugated Er-Te slabs and Te square nets. (b) Schematic of the Fermi surface showing the nesting vectors (adapted from Ref. [43]). (c) Scattering geometry of the experiment with reference to the sample orientation. ki and ks represent momenta of the incident and scattered electron, respectively, and $q$ is the in-plane momentum transfer which is along $c$-axis.
@@
Figure 2: Elastic response: (a) M-EELS spectra ($\omega\approx 0$) as a function of $q$ for different temperatures. (b) The integrated area under the CDW peak $q\approx q_{0}$ obtained from fits to (2) along with the fit to mean-field interpolation formula, normalized to the value in the $T\rightarrow 0$. (c) $w(t)=\gamma_{\text{CDW}}/\gamma_{(0,1)}$, the ratio of the width of the CDW satellite to that of the Bragg peak as a function of temperature. (d) Ordering wave vector $q_{0}$ as a function of temperature. Vertical gray dashed line in (c,d) denotes the Tc obtained in fit in (b).
@@
where $G(Q,T)=1/[\kappa Q^{2}+\mu(T)]$ is a Lorentzian, $\sigma$ represents the disorder strength, $Q=q-q_{0}$ is the momentum relative to the CDW wave vector, and the other parameters quantify the CDW stiffness and correlation length. The first term in (1) is truly static, while the second is dynamical and, in principle, could be eliminated by using an energy analyzer that perfectly isolated the elastic ($\omega=0$) response.
@@
where $q_{0}(T)$ is the CDW ordering wavevector and the width, $\gamma_{q}(T)$, is proportional to the inverse correlation length (see Supplemental Information). The sum $m(T)q+c(T)$ represents a linear background whose meaning will be discussed further below. Surprisingly, adding an additional, Lorentzian-squared term did not improve the quality of the fits. This is surprising given that our measurement is energy-resolved, with an overall resolution of 5.6 meV, and should isolate the Lorentzian-squared term in (1). We conclude that the CDW in ErTe3 is influenced by disorder, but extremely weakly.
@@
Figure 3: Dynamical response: Low energy M-EELS spectra normalized to the maximum intensity for different temperatures for (a) $q=0.6r.l.u.$ and (b) $q\sim q_{0}$. (c) $\gamma(q,T)$ extracted by fitting the loss spectra to (4). The dashed lines are a guide to the eye. (d) Temperature dependence of $\gamma(q,T)$ for $q=0.6$ (r.l.u.) and $q\approx q_{0}$. (e) $\tau^{-1}$ and (f) $D$ obtained from fitting $\gamma(q,T)$ near $q_{0}$ to (5).
@@
The integrated intensity of elastic CDW reflection, $I_{\text{CDW}}(T)$, is plotted in Fig.2(b). $I_{\text{CDW}}(T)$ is proportional to the square of the order parameter, $\Delta(T)$, and a fit to the BCS interpolation formula $\Delta(T)=\Delta_{0}\tanh\left({1.74\sqrt{T_{c}/T-1}}\right)$, yields a $T_{C}=268(7)K$ which is consistent with the previous x-ray diffraction measurements [19]. In the intermediate regime between 100-250K, the integrated intensity is smaller than the mean field prediction, which was also observed in the evolution of the single particle gap in ARPES [20]. Vestiges of the CDW order is also observed above $T_{c}$, up to $T=$290 K, presumably due to pinning on disorder [5].
@@
Fig.2(c) shows the ratio of the width of the CDW to that of the (0,1) Bragg peak, i.e., $w(t)=\gamma_{\text{CDW}}/\gamma_{(0,1)}$, as a function of temperature. Below, 225K the width of the CDW is nearly the same as that of the Bragg peak, indicating that it is limited by the instrumental resolution. This further supports the conclusion that disorder plays little role in the properties of this CDW. Above 225K, the CDW peak broadens, indicating that the CDW correlation length gets shorter as the temperature is increased through $T_{c}$. The CDW wavevector, $q_{0}(T)$, shows a non-monotonic dependence on temperature (Fig.2(d)). This behavior has also been observed in x-ray measurements of other tritellurides [19] and is a common characteristic of incommensurate CDWs. The exact value of $q_{0}$ often reflects a balance between factors such as the Fermi surface morphology, lattice pinning, the compressibility of uncondensed electrons, and coupling to hidden order parameters [50, 51, 49]. These behaviors of $q_{0}(T)$ suggests that the CDW in ErTe3 is shaped by a combination of these effects.
@@
Having established the structural properties of the CDW through elastic scattering, we now focus on its collective dynamics. Figures 3(a,b) present the temperature dependence of the M-EELS spectra at $q=0.6$ r.l.u. and near the CDW ordering wavevector, $q=q_{0}$. Across all temperatures and momenta, the spectra are dominated by a quasi-elastic response, with no evidence of dispersive collective excitations. This finding contrasts sharply with IXS studies, which reveal a softening of phonon modes near the CDW wavevector above $T_{C1}$ [25, 26]. Since M-EELS is more sensitive to valence band electronic excitations, while IXS primarily detects phonons, this suggests that the electronic excitations in ErTe3 behave quite differently from its lattice excitations. This conclusion is further reinforced by the absence of phonon peaks in the low-energy optical conductivity of ErTe3 [22], an effect attributed to strong metallic screening.
@@
Despite the absence of well-defined, propagating collective modes, the dynamical charge susceptibility shows a clear temperature dependence. A broadening of the quasi-elastic response is evident in the raw spectrum (Fig.3(a,b)), suggesting that the CDW undergoes nontrivial, relaxational dynamics at an energy scale of $\sim k_{B}T$.
@@
where $n(\omega,T)$ is the Bose occupation factor and $V_{\text{eff}}(\omega,q)$ is the Coulomb matrix element [32]. This minimal model, with only two adjustable parameters ($A(q,T)$ and $\gamma(q,T)$) provides excellent fits to the M-EELS data (see Supplemental Information), confirming that the charge dynamics in ErTe3 are predominantly relaxational.
@@
$\gamma(q,T)=\hbar\tau^{-1}+\hbar D(T)\;(q-q_{0})^{2}$ (5)
where $\tau^{-1}$ represents pure dissipation and $D(T)$ is the diffusion constant. Figs. 3(e,f) show plots of $\tau^{-1}$ and $D(T)$, respectively. The dissipation rate, $\tau^{-1}$, varies quadratically with temperature, consistent with the behavior of a good Fermi liquid. Intriguingly, $D(T)$ peaks around $\sim$250 K, slightly below $T_{c}$ where a strong violation of the Wiedemann-Franz law was previously observed in thermal transport measurements [53]. This suggests that electronic CDW excitations propagate diffusively, with the diffusivity reaching its maximum just below $T_{c}$, when the CDW amplitude is large but still subject to significant thermal fluctuations. The characteristic diffusion length, $\lambda$, defined as $\lambda\sim\sqrt{D\;\tau}$, ranges between $\sim$ 2–10 lattice constants, comparable to that observed in the “stripe-ordered” phase La2-xBaxCuO4 $\left(x\sim\frac{1}{8}\right)$ with time-resolved x-ray scattering experiments [54].
@@
Finally, our analysis allows us to quantify the behavior of the static susceptibility, $\chi(q)$, which is traditionally expected to diverge near a CDW transition [1]. The static susceptibility is given by the real part of the complex susceptibility $\chi(q,\omega)$ evaluated at zero frequency. While this could be obtained via a Kramers-Kronig transformation of the data, we can also directly compute it from the real part associated with (3),
@@
Finally, we note that, while the largest changes in the static susceptibility occur at $q_{0}$, the susceptibility shows significant temperature dependence at all momenta measured, which appear as a linear background in (2). This indicates that a substantial fraction of the valence electrons—even those not directly involved in CDW ordering—contribute significantly to the overall change in charge susceptibility. What underlying physics would drive such fluctuations everywhere in momentum space is an open question.
@@ Acknowledgements.
We thank Steven A. Kivelson and Anisha G. Singh for helpful discussions. M-EELS measurements were supported by the Center for Quantum Sensing and Quantum Materials, an Energy Frontier Research Center funded by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award DE-SC0021238. Q.J. and I.R.F (crystal growth and characterization) were supported by DOE BES contract DE-AC02-76SF00515. T.C.C (surface preparation) was supported by DOE BES contract DE-FG02-07ER46383. E.F. (theory) was supported by National Science Foundation grant DMR 2225920. Q.J. was also partially supported by a Geballe Laboratory for Advanced Materials Postdoctoral Fellowship. P.A. gratefully acknowledges support from the EPiQS program of the Gordon and Betty Moore Foundation, grant GBMF9452.